OpenStudy (anonymous):

An airplane flew 3 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. 380 mph 340 mph 320 mph 300 mph

OpenStudy (anonymous):

Let x=Speed of plane in still air y=Speed of head wind y=60 3(x-60)=2(x+60) 3x-180=2x+120 x=120+180 x=300

OpenStudy (anonymous):

thank you